The sum of two numbers is 28. One number is 3 times as big as the other. What is the smaller number?
Answer by ewatrrr(24785)


Hi
sum of two numbers is 28
Let x and (28-x) represent the two numbers
question states***
x = 3(28-x)
solving for x
x = 84 - 3x
4x = 84
x = 21 and 7 (28-21) are the two numbers. 7 the smaller.
CHECKING our Answer***
21 = 3*7